Consider the following numerical example of the simple Keynesian model with no government spending, taxes, or a foreign sector (all figures in R millions):
C = 100 + 0,9Y
I = 50
Answer the following questions.

(a) Use a graph to illustrate the equilibrium level of output.5

(b) Suppose the level of output that creates full employment in the economy is 1 800. Using the multiplier, determine the level of investment spending that would create full employment in this economy.5
1
Expert's answer
2020-11-27T13:36:09-0500

a)

The equilibrium level of output is:

Y = C+ I = 100 + 0.9Y + 50,

0.1Y= 150,

Y =1500.

The equilibrium level of output is shown as an intersection point of AD and AS(Y) curves.


b)

Our new Output value Y = 1800

Y = C+I

1800 = 100 + (0.9x1800) + I

I = 1800 -100 - 1620

Investment Spending (I) = 80